# AirRAG: Activating Intrinsic Reasoning for Retrieval Augmented Generation using Tree-based Search

### v2

@@ 4.1.2 Implementation Details
Table 1: Overall evaluation results on the test sets of three datasets. * indicates the results reproduced by us. $L_{\max}$ denotes the maximum number of input tokens across all rollouts. The best results for each $L_{\max}$ are in bold. The number of both rollouts and output sequences is set to 1.
4.1.3 Baselines and Metrics
To investigate the enhancement effects of thinking and planning on complex RAG tasks, we compare it with vanilla RAG, which performs only a single retrieval and generation process. We evaluate the naive generators of Qwen2.5-14B-Instruct (Yang et al., 2024) and Llama-3-8B-Instruct (Grattafiori et al., 2024). In the retrieval phase, we employ multilingual-e5-base (Wang et al., 2024) as the retriever. The prompt of vanilla RAG are shown in the Appendix C. For iterative retrieval, we compare AirRAG with Iter-RetGen (Shao et al., 2023), Self-RAG (Asai et al., 2024), Auto-RAG (Yu et al., 2024), and IterDRAG (Yue et al., 2024). To further explore RAG performance and inference computation scaling, we focus on a comparison with IterDRAG for a given budget on inference computation. For evaluation metrics, we report Exact Match (EM), F1 score (F1) and Accuracy (Acc) between the generated summary and gold answer, where accuracy measures whether the gold answer is covered in the generated answer.
4.2 Main Results (RQ1)
We first evaluate the performance of AirRAG on various complex QA datasets. Table 1 compares its accuracy and F1 with strong baselines under the given inference computation budget, which is implemented based on Qwen2.5-14B-Instruct and one million document database. The optimal performance exhibits consistent gains as $L_{\max}$ expands, which is termed as the inference scaling laws for RAG (Yue et al., 2024). We integrate the remaining methods for a given maximum computational budget into our approach, dubbed as AirRAG-Blender. The best results are obtained by using only the SA action to refine the final answer from all candidates, as shown in Table 1. This also demonstrates the flexibility of our approach architecture. In addition, to verify the robustness and generalization of AirRAG, Table 5 shows the performance on more diverse LLMs and datasets. For a fair comparison, we utilize the widely used Wikipedia dump from December 2018 (Karpukhin et al., 2020) as the retrieval corpus. We observe consistent improvements over vanilla RAG and existing iterative methods (more than 10% on average). The significant boost over IterDRAG and Auto-RAG suggests that AirRAG explores more effective reasoning paths through the human-like thinking paradigm and tree-based search.
Method Average
F1 Acc
Vanilla RAG 47.0 43.2
IterDRAG 49.8 45.9
AirRAG
+ $n_{\text{all}}$=1 62.9 58.8
+ $n_{\text{all}}$=3 63.4 62.1
[…]
+ $n_{a_{1},a_{4}}$=3, $n_{a_{2},a_{3},a_{5}}$=1, $q_{div}$=1.0 65.1 63.9
Table 2: Performance comparison with different computationally optimal strategies. $n_{a_{i}}$ denotes the number of output sequences of the reasoning action $a_{i}$ in a single extension. $q_{div}$ indicates that setting top-$p$ to 1.0 and temperature to 1.0 for query-related actions, i.e. SAY and QT, increases the diversity of reasoning. The default sampling parameters top-$p$, top-$k$ and temperature are set to 0.8, 50 and 0.7 respectively. Rational sampling strategies and inference scaling further improve performance across multiple datasets.
4.3 Inference Scaling for RAG (RQ2)
Inference computation scaling can enable LLMs to improve their output performance (Snell et al., 2024). Self-consistency can also improve the robustness of the reasoning process (Wang et al., 2023). Therefore, we carry out a comprehensive experimental analysis on the inference computation scaling. Based on tree-based search and RAG scenario, there are multiple ways to optimize the use of inference computation resources. Specifically, we can adjust both the number of retrieved documents in a single retrieval and the effective context length in all iterations. The average performance of three datasets exhibits consistent gains in Figure 3. In subsequent experiments, unless otherwise specified, the data presented represent the average performance across the HotpotQA, MuSiQue, and 2Wiki datasets. In particular, the initial computation scaling brings significant performance improvements. In addition, the number of output sequences and rollouts in MCTS can expand the solution space and explore potential reasoning paths. As shown in Figure 1, the average performance increases with the number of output sequences per action, demonstrating the effectiveness of self-consistency. We also investigate the number of effective reasoning paths under different rollouts in Figure 4. Similarly, the performance improvement caused by the increase of effective reasoning paths in the early stage is relatively high. Similarly, the increase in early reasoning paths leads to relatively higher performance gains. We provide additional dataset-specific results in Appendix B.
4.4 Ablation Studies
Effect of Computationally Optimal Strategies (RQ3). Extensive experiments show that the outputs of certain actions (e.g., RA, DA and SA) are almost consistent when performing multiple generations. Therefore, we only increase the number of output sequences (short for $n$) for the remaining actions (e.g., SAY and QT), which reduces invalid inference computation while maintaining good results. This also reflects that this kind of reasoning action, which effectively activates the creativity of LLMs, requires more diversified sampling strategies. We adjust the sampling parameters (top-$p$=1.0 and temperature=1.0) to improve the diversity of the model output. The complete experimental results in Table 2 show that the diversity of key actions can significantly improve performance.
4.5 Qualitative Analysis (RQ5)
To make it easier to understand why our proposed AirRAG works, we present a qualitative analysis in MuSiQue. Existing iterative methods are often trapped in a single solution space when confronted with complex tasks. As illustrated in Figure 13, these iterative methods exhibit a key limitation that insufficient or ambiguous retrieval context can lead to repetitive follow-up queries until it reaches the predefined maximum depth of iterations. This inefficient iteration results in high computational cost and incorrect answer. In contrast, our proposed AirRAG designs efficient reasoning actions to activate the intrinsic reasoning capabilities of LLMs. As shown in Figure 14, the SAY action decomposes the original query into a more rational sequence of sub-queries, and then the combination of RA and QT ensures the accuracy of the intermediate reasoning step. We eventually leverage the efficient reasoning trajectory to obtain the correct answer.
5 Conclusions
In this paper, we propose AirRAG, a novel RAG approach to activate intrinsic reasoning capabilities of LLMs. AirRAG designs an efficient action space for the controllable reasoning generation. We also introduce Monte Carlo Tree Search to expand the solution space. Meanwhile, by employing the tree-based search and self-consistency verification, we explore potential reasoning paths and achieve comprehensive inference computation scaling in RAG. In addition, computationally optimal strategies are used to apply more computation to key actions, leading to further performance improvements. Experimental results on diverse QA datasets demonstrate the significant superiority of AirRAG over other methods designed for complex RAG scenarios. Furthermore, AirRAG can be integrated with other advanced methods with great flexibility.
Limitations
Although our model achieves competitive performance in various RAG tasks, there are some methods and limitations that can be improved. The current optimal computation allocation strategy is derived from sufficient experiments. We can consider designing an automated policy model to implement the trade-off between computational cost and performance. Despite great efforts in the inference scaling of RAG, the experimental analysis may be limited due to the massive computational cost of tree-based search approaches. We will explore more complex reasoning tasks to verify the robustness and effectiveness of our approach. In addition, the large search space also brings more noise information, so we will further investigate the reward model or strategy to explore a better reasoning path.
Appendix A Implementation Details
In the retrieval process, we employ the multilingual-e5-base (Wang et al., 2024) as the retriever and use the widely used Wikipedia dump from December 2018 as the retrieval corpus (Karpukhin et al., 2020) which comprises over 21 million passages. For generation, the default sampling parameters top-$p$, top-$k$ and temperature are set to 0.8, 50 and 0.7 respectively. Evaluation metrics include Exact Match (EM), F1 score (F1), and Accuracy (Acc), where accuracy indicates whether the ground truth is a substring of the final generated answer. For reward model training, we sample 8,000 question-answer pairs from each dataset and generate more than 156,000 reasoning paths using our proposed AirRAG (rollouts=32, $n$=4, $q_{div}$=1.0). In inference scaling experiments, we sample maximum computation budgets $L_{\max}$ (e.g., 8k, 16k, 32k, 64k and 128k tokens). The $L_{\max}$ (maximum effective context length) denotes the maximum number of input tokens across all rollouts following  (Yue et al., 2024). The predetermined maximum tree depth $d$ is set to 10, specifically indicating that the SAY and SA actions are executed once, while the RA-QT or QT-RA actions have a maximum of 4 iterations.
Appendix B Additional Experiment Results
We report the average performance of our approach on three datasets with the support of the Qwen2.5-72B-Instruct model in Table 3 and 4. The experimental results show that the performance is further improved (+4.6%, +6.1%, +6.9% in average accuracy) over 14B model. Meanwhile, we observe that the performance of our AirRAG based on the 14B model surpasses other methods using a 72B model. Furthermore, we present detailed inference scaling results for each dataset individually, as shown in Figure 7 and Figure 8.
Method Qwen2.5-72B Qwen2.5-14B
F1 Acc F1 Acc
Vanilla RAG 51.6 46.9 47.0 43.2
IterDRAG 53.4 48.5 49.8 45.9
AirRAG 69.6 68.5 65.1 63.9
Table 3: Performance comparison between Qwen2.5-14B-Instruct and Qwen2.5-72B-Instruct. Three documents are retrieved in a single RA action. Rollouts and $n$ are set to 32 and 3 respectively.
$L_{\max}$ Method Qwen2.5-72B Qwen2.5-14B
F1 Acc F1 Acc
8k ZeroShot QA 38.8 37.6 34.7 33.6
Vanilla RAG 56.2 51.8 49.7 45.5
IterDRAG 57.0 52.2 53.8 48.7
AirRAG-Lite 70.9 66.6 63.8 59.1
AirRAG 72.4 67.6 65.6 61.5
AirRAG-Blender 73.2 70.0 68.3 66.0
16k Vanilla RAG 56.8 52.5 53.5 49.2
IterDRAG 58.3 53.9 55.3 50.2
AirRAG-Lite 73.2 68.8 65.2 60.7
AirRAG 74.2 69.7 67.2 62.8
AirRAG-Blender 75.3 71.0 69.3 66.5
Table 4: Overall average evaluation results between Qwen2.5-14B-Instruct and Qwen2.5-72B-Instruct over three datasets. $L_{\max}$ denotes the maximum number of input tokens across all rollouts or iterations. Both rollouts and $n$ are set to 1.
Appendix C Prompt Examples

### v3

@@ 4.1.2 Implementation Details
4.1.3 Baselines and Metrics
To investigate the enhancement effects of thinking and planning on complex RAG tasks, we compare it with vanilla RAG, which performs only a single retrieval and generation process. We evaluate the naive generators of Qwen2.5 (Yang et al., 2024) series instruction models and Llama3-8B-Instruct (Grattafiori et al., 2024). In the retrieval phase, we employ multilingual-e5-base (Wang et al., 2024) as the retriever and utilize the widely used Wikipedia dump from December 2018 as the retrieval corpus (Karpukhin et al., 2020). The prompt of vanilla RAG are shown in the Appendix C. For iterative retrieval, we compare AirRAG with Iter-RetGen (Shao et al., 2023), Self-RAG (Asai et al., 2024), Auto-RAG (Yu et al., 2024), and IterDRAG (Yue et al., 2024). For agentic retrieval, we compare AirRAG with Search-o1 (Li et al., 2025), ReSearch (Chen et al., 2025), and DeepResearcher (Zheng et al., 2025). To further explore RAG performance and inference computation scaling, we focus on a comparison with IterDRAG for a given budget on inference computation. For evaluation metrics, we report Exact Match (EM), F1 score (F1) and Accuracy (Acc) between the generated summary and gold answer, where accuracy measures whether the gold answer is covered in the generated answer.
4.2 Main Results (RQ1)
We first evaluate the performance of AirRAG on various complex QA datasets. Table 1 compares its accuracy and F1 scores with strong baselines based on LLMs of different scales. The optimal performance exhibits consistent gains as the LLMs scale up. For the Qwen2.5-7b-instruct model, our approach achieves the best performance, even surpassing the trainable approaches. To further validate its effectiveness on large-scale reasoning models, we also conduct experiments on Qwen3-235B in both the thinking mode and non-thinking mode. In thinking mode, our approach achieves state-of-the-art performance among all datasets. In addition, to verify the robustness and generalization of AirRAG, Table 5 shows the performance on more diverse LLMs and datasets. We observe consistent improvements over vanilla RAG and existing iterative methods (more than 10% on average). The significant boost over IterDRAG and Auto-RAG suggests that AirRAG explores more effective reasoning paths through the human-like thinking paradigm and tree-based search.
Method NQ TriviaQA HotpotQA MuSiQue 2Wiki Average
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Qwen2.5-7B
ZeroShot QA 38.4 37.4 57.7 56.3 36.1 34.8 9.1 7.5 45.0 44.1 37.3 36.0
Vanilla RAG 57.8 53.9 70.1 66.3 61.3 56.9 13.5 8.3 45.9 42.8 49.7 45.6
IterDRAG 58.3 54.1 73.5 69.1 65.3 60.7 18.3 13.0 51.8 47.0 53.4 48.8
Search-o1 57.8 54.3 72.6 69.8 57.3 54.2 20.2 18.5 56.9 53.6 53.0 50.1
ReSearch 61.3 59.6 76.2 73.4 70.3 68.1 30.9 27.4 62.5 61.0 60.2 57.9
DeepResearcher 62.4 59.8 75.9 73.2 67.6 63.9 33.9 27.8 65.4 62.3 61.0 57.4
AirRAG-Lite 60.8 57.3 74.1 70.0 68.2 63.2 23.4 17.3 51.1 48.0 55.5 51.2
AirRAG 60.4 56.2 74.3 70.1 74.9 70.0 30.3 23.6 65.7 63.2 61.1 56.6
Qwen2.5-14B
ZeroShot QA 47.1 46.3 70.7 69.5 42.5 41.3 13.5 12.1 48.2 47.3 44.4 43.3
Vanilla RAG 62.0 58.2 75.2 71.2 67.0 62.0 20.0 14.4 52.0 49.4 55.2 51.0
IterDRAG 58.5 53.9 76.4 72.3 71.8 66.2 23.9 17.2 57.4 54.1 57.6 52.7
Search-o1 61.2 59.7 74.3 72.6 71.4 67.4 23.2 20.4 58.1 55.8 57.6 55.2
AirRAG-Lite 64.8 60.9 78.9 75.0 77.2 72.5 30.3 24.3 70.3 68.0 64.3 60.1
AirRAG 66.2 62.1 78.1 73.8 79.9 75.3 36.0 31.9 70.4 68.7 66.1 62.4
Qwen2.5-32B
ZeroShot QA 46.8 45.9 69.8 68.8 42.9 41.8 11.1 9.6 48.1 47.2 43.7 42.7
Vanilla RAG 60.6 56.7 75.4 71.6 66.5 62.5 19.5 13.8 51.8 49.6 54.7 50.8
IterDRAG 61.1 56.7 76.9 73.0 72.0 67.3 23.3 17.9 58.2 55.4 58.3 54.1
Search-o1 63.5 61.6 75.6 72.8 72.8 68.4 30.2 27.3 60.5 58.9 60.5 57.8
ReSearch 63.8 61.2 74.6 72.3 76.2 72.6 38.3 33.4 66.8 62.8 63.9 60.5
AirRAG-Lite 65.6 61.9 75.8 71.6 79.3 74.8 33.3 32.5 72.0 70.8 65.2 62.3
AirRAG 66.5 62.7 78.9 74.6 81.1 76.1 36.5 32.7 71.9 70.6 67.0 63.3
Qwen3-235B (non-thinking)
ZeroShot QA 64.1 63.7 77.1 76.3 53.9 52.9 17.1 15.7 56.8 56.1 53.8 52.9
Vanilla RAG 66.0 65.3 78.0 76.2 69.2 67.7 20.8 19.1 53.3 52.2 57.4 56.1
IterDRAG 65.7 63.3 78.8 76.9 75.5 69.7 32.2 25.2 64.8 60.8 63.4 59.2
Search-o1 67.3 66.2 77.3 76.4 75.9 73.8 36.7 34.2 71.3 68.2 65.7 63.7
AirRAG-Lite 67.7 66.9 77.6 75.8 78.3 76.8 43.7 36.5 74.9 74.1 68.4 65.7
AirRAG 66.4 65.6 79.1 77.3 79.6 78.1 47.2 40.0 76.2 75.5 69.7 67.3
Qwen3-235B (thinking)
ZeroShot QA 66.1 65.6 79.3 78.6 54.2 53.5 21.3 19.9 56.6 56.3 55.5 54.8
Vanilla RAG 67.6 67.1 77.9 76.9 72.5 71.9 18.2 16.4 57.5 56.6 58.7 57.8
IterDRAG 68.8 67.1 80.9 78.7 76.3 71.1 30.0 23.1 67.1 64.5 64.6 60.9
Search-o1 67.2 66.4 78.1 77.5 75.2 72.4 33.4 28.9 69.2 66.3 64.6 62.3
AirRAG-Lite 73.2 72.7 81.1 80.1 86.2 85.6 44.2 37.8 76.4 75.6 72.2 70.3
AirRAG 74.3 72.8 81.4 80.1 84.7 84.0 47.5 40.3 76.8 76.2 72.9 70.7
Table 1: Overall evaluation results on the test sets of five datasets. * indicates the results reproduced by us. The best results for each model are in bold. The number of both rollouts and output sequences is set to 1. The number of documents for a single retrieval is set to 5.
4.3 Inference Scaling for RAG (RQ2)
Inference computation scaling can enable LLMs to improve their output performance (Snell et al., 2024). Self-consistency can also improve the robustness of the reasoning process (Wang et al., 2023). Therefore, we carry out a comprehensive experimental analysis on the inference computation scaling. Based on tree-based search and RAG scenario, there are multiple ways to optimize the use of inference computation resources. Specifically, we can adjust both the number of retrieved documents in a single retrieval and the effective context length in all iterations. The average performance of three datasets exhibits consistent gains in Figure 3. In subsequent experiments, unless otherwise specified, the data presented represent the average performance across the HotpotQA, MuSiQue, and 2Wiki datasets. In particular, the initial computation scaling brings significant performance improvements. In addition, the number of output sequences and rollouts in MCTS can expand the solution space and explore potential reasoning paths. As shown in Figure 1, the average performance increases with the number of output sequences per action, demonstrating the effectiveness of self-consistency. We also investigate the number of effective reasoning paths under different rollouts in Figure 4. The performance improvement caused by the increase of effective reasoning paths in the early stage is relatively high. We provide additional dataset-specific results in Appendix B.
Method Average
F1 Acc
Vanilla RAG 47.0 43.2
IterDRAG 49.8 45.9
AirRAG
+ $n_{\text{all}}$=1 62.9 58.8
+ $n_{\text{all}}$=3 63.4 62.1
[…]
+ $n_{a_{1},a_{4}}$=3, $n_{a_{2},a_{3},a_{5}}$=1, $q_{div}$=1.0 65.1 63.9
Table 2: Performance comparison with different computationally optimal strategies on the HotpotQA, MuSiQue and 2Wiki datasets. $n_{a_{i}}$ denotes the number of output sequences of the action $a_{i}$ in a single extension. $q_{div}$ indicates that setting top-$p$ to 1.0 and temperature to 1.0 for query-related actions, i.e. SAY and QT, increases the diversity of reasoning. The default sampling parameters top-$p$, top-$k$ and temperature are set to 0.8, 50 and 0.7 respectively. Rational sampling strategies further improve performance across multiple datasets.
4.4 Ablation Studies
Effect of Computationally Optimal Strategies (RQ3). Extensive experiments show that the outputs of certain actions (e.g., RA, DA and SA) are almost consistent when performing multiple generations. Therefore, we only increase the number of output sequences (short for $n$) for the remaining actions (e.g., SAY and QT), which reduces invalid inference computation while maintaining good results. This also reflects that this kind of reasoning action, which effectively activates the creativity of LLMs, requires more diversified sampling strategies. We adjust the sampling parameters (top-pp=1.0 and temperature=1.0) to improve the diversity of the model output. The complete experimental results in Table 2 show that the diversity of key actions can significantly improve performance.
$L_{\max}$ Method database_size retrieval_time retrieval_number e2e
8k Vanilla RAG 100w 0.900 1.00 3.828
IterDRAG 100w 1.833 2.04 6.703
AirRAG-Lite 100w 2.205 2.25 8.482
AirRAG 100w 3.453 3.84 12.752
Table 3: Performance analysis of inference efficiency. $L_{\max}$ denotes the maximum number of input tokens across all rollouts. The retrieved database contains approximately one million documents. e2e and retrieval_time denote the average total time for a single question-answering process and the time spent on retrieval respectively, measured in seconds. Other inference configurations remain consistent with those in Table 4.
4.5 Inference Efficiency and Qualitative Analysis (RQ5)
Given the inherently large search space for the tree-based search, we design computational optimization strategies for different actions to avoid redundant and inefficient expansions, as shown in Table 2. Furthermore, in Section 3.2.2, we propose pruning strategies for state nodes and reasoning paths. These optimizations significantly reduce inefficient LLM inference and repetitive path exploration. In practical applications, we can select appropriate configuration parameters such as rollout, $n$, and $L_{\max}$ based on computational resource budgets and time constraints, ensuring effectiveness while achieving inference efficiency comparable to current mainstream iterative RAG approaches. We analyze the average inference efficiency per sample on the HotpotQA dataset in Table 3. By combining these results with those reported in Table 4 and Table 1, we are able to analyze the trade-offs between computational cost and performance.
To make it easier to understand why our proposed AirRAG works, we present a qualitative analysis in MuSiQue. Existing iterative methods are often trapped in a single solution space when confronted with complex tasks. As illustrated in Figure 13, these iterative methods exhibit a key limitation that insufficient or ambiguous retrieval context can lead to repetitive follow-up queries until it reaches the predefined maximum depth of iterations. This inefficient iteration results in high computational cost and incorrect answer. In contrast, AirRAG designs efficient reasoning actions to achieve autonomous planning and reasoning. As shown in Figure 14, the SAY action decomposes the original query into a more rational sequence of sub-queries, and then the combination of RA and QT ensures the accuracy of the intermediate reasoning step. We eventually leverage the efficient reasoning trajectory to obtain the correct answer.
5 Conclusions
In this paper, we propose AirRAG, a novel RAG approach to fully leverage the planning and reasoning capabilities of LLMs. AirRAG designs an efficient action space for the controllable reasoning generation. We also introduce Monte Carlo Tree Search to expand the solution space. Meanwhile, by employing the tree-based search and self-consistency verification, we explore potential reasoning paths and achieve comprehensive inference computation scaling. In addition, computationally optimal strategies are used to apply more computation to key actions, leading to further performance improvements. Experimental results on diverse QA datasets demonstrate the significant superiority of AirRAG over other methods designed for complex deep search scenarios.
Limitations
Although our model achieves competitive performance in various RAG tasks, there are some limitations that can be improved. The current optimal computation allocation strategy is derived from sufficient experiments. We can consider designing an automated policy model to implement the trade-off between computational cost and performance. Despite great efforts in the inference scaling of RAG, the experimental analysis may be limited due to the massive computational cost of tree-based search approaches. We will explore more complex reasoning tasks to verify the robustness and effectiveness of our approach. In addition, the large search space also brings more noise information, so we will further investigate the reward model or strategy to explore a better reasoning path.
Appendix A Implementation Details
In the retrieval process, we employ the multilingual-e5-base (Wang et al., 2024) as the retriever and use the widely used Wikipedia dump from December 2018 as the retrieval corpus (Karpukhin et al., 2020) which comprises over 21 million passages. For generation, the default sampling parameters top-$p$, top-$k$ and temperature are set to 0.8, 50 and 0.7 respectively. Evaluation metrics include Exact Match (EM), F1 score (F1), and Accuracy (Acc), where accuracy indicates whether the ground truth is a substring of the final generated answer. For reward model training, we sample 8,000 question-answer pairs from each dataset and generate more than 156,000 reasoning paths using our proposed AirRAG (rollouts=32, nn=4, $q_{div}$=1.0). In inference scaling experiments, we sample maximum computation budgets $L_{\max}$ (e.g., 8k, 16k, 32k, 64k and 128k tokens). The $L_{\max}$ (maximum effective context length) denotes the maximum number of input tokens across all rollouts following  (Yue et al., 2024). The predetermined maximum tree depth $d$ is set to 10, specifically indicating that the SAY and SA actions are executed once, while the RA-QT or QT-RA actions have a maximum of 4 iterations.
To further substantiate our experimental approach, we present comprehensive details and rationales for the experimental design, including the parameter configurations for the rollout process in MCTS, the selection of the UCT function, reward computation strategies, and the implementation of baseline methods.
Rollout Setting for MCTS. In some experimental configurations, we adopt a rollout setting of $1$ for MCTS. A rollout value of $1$ does not imply that MCTS is not executed, but rather that only one complete simulation is performed at each decision step to estimate the value of the current node. This setting reduces computational cost and allows for rapid generation of preliminary results, which is especially beneficial in resource-constrained scenarios. Figure 4 presents the performance across rollout values ranging from $1$ to $32$, while Table 2 reports results obtained with a rollout value of $32$.
Contribution of MCTS. MCTS primarily operates on actions such as QT, RA, and SA, as well as parameters (e.g., $n$, $\mathrm{top}_{p}$, $\mathrm{top}_{k}$) within each action. As illustrated in Figure 4 and Table 2, applying MCTS with higher rollout values significantly improves performance compared to using a rollout value of $1$, and outperforms baselines such as IterDRAG and Search-o1.
UCT and Implicit Priors. We explore both UCT and PUCT as potential search strategies for MCTS and ultimately adopt UCT in our final implementation, owing to its simplicity and consistent empirical performance. Although UCT does not explicitly incorporate prior probabilities as in PUCT, our framework implicitly introduces a form of policy prior through dependencies between actions, which constrain action transitions. This implicit prior serves a function analogous to the explicit $P(s,a)$ calculation in PUCT, guiding the search toward high-quality actions based on prior reasoning steps. The complete action execution process is detailed in Section 3.1.
Reward Computation and Backpropagation. The reward score is determined by whether the final correct answer is obtained. Specifically, after executing action $A_{5}$ (Summary-Answer), a reward score is computed for each leaf node. In our experimental setup, we consider scenarios in which the self-consistency sampling count is set to $1$ or $3$, as summarized in Table 1 with $n_{\text{all}}=1/3$. When $n_{\text{all}}=1$, the $A_{5}$ action generates a single answer and the leaf node is assigned a reward $Q$ of 1. For $n_{\text{all}}=3$, we apply the clustering method (e.g., jcdScore) described in Section 3.3 to calculate reward scores for the three answer nodes. This method assigns higher scores to answers with greater confidence, consistent with the voting strategies commonly employed in mathematical tasks. We then use these computed scores to backpropagate and update the scores for all nodes along the search path. In the formula $Q(s_{i},a_{i})=Q(s_{i},a_{i})+Q(s_{d},a_{d})$, $Q(s_{d},a_{d})$ denotes the reward value associated with the final answer.
Self-Consistency Baselines. For the Vanilla RAG method, the improvement brought by self-consistency voting is limited. In our main experiments, we set the rollout parameter to $1$ to ensure a fair comparison and to avoid potential confounding effects arising from applying self-consistency solely to the final results.
Context Length Expansion for IterDRAG. For IterDRAG, we follow the protocol outlined in the original paper, modifying three key aspects: the number of retrieved documents (ranging from $3$ up to $300$ maximum), the number of contextual examples (fixed at $5$ in our experiments), and the number of iterations (up to $5$). Further details regarding these inference scaling strategies are discussed in the work by Google DeepMind (Yue et al., 2024).
Appendix B Additional Experiment Results
We evaluate the performance of AirRAG on various complex QA datasets. Table 4 compares its accuracy and F1 with strong baselines under the given inference computation budget, which is implemented based on Qwen2.5-14B-Instruct and one million document database. The optimal performance exhibits consistent gains as $L_{\max}$ expands, which is termed as the inference scaling laws for RAG (Yue et al., 2024). We integrate the remaining methods for a given maximum computational budget into our approach, dubbed as AirRAG-Blender. The best results are obtained by using only the SA action to refine the final answer from all candidates, as shown in Table 4. This also demonstrates the flexibility of our approach architecture. In addition, to verify the robustness and generalization of AirRAG, Table 5 shows the performance on more diverse LLMs and datasets. For a fair comparison, we utilize the widely used Wikipedia dump from December 2018 (Karpukhin et al., 2020) as the retrieval corpus. We observe consistent improvements over vanilla RAG and existing iterative methods (more than 10% on average). The significant boost over IterDRAG and Auto-RAG suggests that AirRAG explores more effective reasoning paths through the human-like thinking paradigm and tree-based search. Furthermore, we present detailed inference scaling results for each dataset individually, as shown in Figure 6 and Figure 8.
@@ Appendix B Additional Experiment Results
Table 4: Overall evaluation results under different computational resource budgets, where Qwen2.5-14B-Instruct is used as the generator LLM. * indicates the results reproduced by us. $L_{\max}$ denotes the maximum number of input tokens across all rollouts. The best results for each $L_{\max}$ are in bold. The number of both rollouts and output sequences is set to 1 for our proposed AirRAG methods.